1.    The first term of an arithmetic series is 1 . the common difference of the series is 6.Find the tenth term of the series if the sum of the first n terms of the series is 7400. Show that 3n^2-2n-7400=0


1
Expert's answer
2022-05-04T11:47:37-0400
"a_1=1, d=6"

"a_n=a_1+d(n-1)=1+6(n-1)"

"S_n=\dfrac{a_1+a_n}{2}\cdot n=\dfrac{1+1+6(n-1)}{2}\cdot n"

"=n+3n(n-1)=3n^2-2n"

Given "S_n=7400."


"3n^2-2n=7400"

"3n^2-2n-7400=0, n\ge 1"

"(3n+148)(n-50)=0"

Since "n\ge 0," we take "n=50."


"a_{50}=1+6(50-1)=295"




"a_{10}=1+6(10-1)=55"